# SPOT: Scalable 3D Pre-training via Occupancy Prediction for Learning Transferable 3D Representations

### v3

@@ 1 Introduction
To overcome both task-level and dataset-level gaps and learn general representations, we propose SPOT, namely Scalable Pre-training via Occupancy prediction for learning Transferable representation. Firstly, we argue that occupancy prediction serves as a more general pre-training task for task-level generalization, as compared to 3D object detection and LiDAR semantic segmentation. The reason lies in that occupancy prediction is based on denser voxel-level labels with abundant classes, which incorporates spatial information similar to 3D object detection as well as semantic information introduced in semantic segmentation. Besides, we consider temporally sufficient representations in the context of 3D pre-training, which contains the information shared among consecutive frames, and theoretically explain why the proposed occupancy-based pre-training outperforms self-supervised methods like MAE in downstream tasks for autonomous driving. Secondly, as the existing datasets use LiDAR sensors with various numbers of laser beams and different category annotation strategies, we propose to use beam re-sampling for point cloud augmentation and class-balancing strategies to overcome these domain gaps. Beam re-sampling augmentation simulates LiDAR sensors with different numbers of laser beams to augment point clouds from a single source pre-training dataset, alleviating the domain gap brought by LiDAR types. Class-balancing strategies apply balance sampling on the dataset and category-specific weights on the loss functions to narrow down the annotation gap. Furthermore, our semi-supervised pre-training experiments as described in Sec. 4.4 provide strong evidence that SPOT consistently enhances the performance of different downstream tasks, even when only a minimal number of labels are used. Last but not least, we observe that using larger amounts of pre-training data leads to better performance on various downstream tasks. This holds true even when the pre-training data are generated through pseudo-labeling. These findings indicate that SPOT is a scalable pre-training method for LiDAR point clouds, paving the way for large-scale 3D representation learning in autonomous driving.
@@ 3 The Proposed Method
We discuss the proposed SPOT in detail. As shown in Fig. 2, SPOT contains four parts: (a) Augmentations on LiDAR point clouds. (b) Encoder for LiDAR point clouds to generate BEV features, which are pre-trained and used for different downstream architectures and tasks. (c) Decoder to predict occupancy based on BEV features. (d) Loss function with class-balancing strategy. We first introduce the problem formulation as well as the overall pipeline in Sec. 3.1. Then we respectively discuss beam re-sampling augmentation and class-balancing strategies in Sec. 3.2 and Sec. 3.3.
@@ 3.3.1 Dataset Balancing
where $N_{i}$ is the number of samples for the $i^{th}$ class. Fewer samples in a category brings higher weight $s_{i}$ for it. Based on the sampling weights, we can employ a random duplication to balance the classes and compose the final dataset to alleviate the class imbalance. This is advantageous as it allows us to learn scene representations more effectively in the pre-training task, facilitating downstream tasks.
@@ 4 Experiments
The goal of pre-training is to learn general representations for various downstream tasks, datasets, and architectures. We design extensive experiments to answer the question whether SPOT learns such representations in a label-efficiency way. We introduce the employed datasets and experiment setup in Sec. 4.1 and 4.2, respectively, followed by main results with different baselines in Sec. 4.3. Then in Sec. 4.4, we further conduct semi-supervised pre-training experiments specifically to demonstrate the applicability of SPOT in the case of utilizing very small part of annotations, and SPOT consistently demonstrates excellent performance on downstream tasks. Finally, we provide discussions about upstream pre-training and downstream fine-tuning, ablation study and visualization results in Sec. 4.5 and Sec. 4.6.
@@ 4.1.4 ONCE Dataset
ONCE [20] is a large-scale autonomous dataset collected in China using a 40-beam LiDAR. It encompasses a diverse range of data collected at various times, under different weather conditions, and across multiple regions. The dataset comprises over one million frames of point cloud data, with approximately 15K frames containing annotations. The remaining unlabeled point cloud data serves as resources for unsupervised and semi-supervised algorithms.
TABLE I: Few-shot performance of SPOT on nuScenes. P.D.A. denotes Pre-training Data Amount. We fine-tune on 5% nuScenes training data.
@@
TABLE II: Few-shot performance ($\text{AP}_{\text{3D}}$) of SPOT on KITTI benchmark. P.D.A. represents the Pre-training Data Amount, and fine-tuning is performed on 20% KITTI training data.
@@ 4.1.5 SemanticKITTI Dataset
SemanticKITTI dataset [19] is a large-scale dataset based on the KITTI vision benchmark, collected by a 64-beam LiDAR sensor. It has 22 sequences, of which sequences 0-7 and 9-10 are used as the training set (19K frames in total), and sequence 8 (4K frames) is used as the validation set, and the remaining 11 sequences (20K frames) as the test set.
@@ 4.2 Experimental Setup
TABLE III: Few-shot performance of SPOT on SemanticKITTI for segmentation task using 100% pre-training data. We fine-tune on 10% training data and show the results of some of the categories.
@@ 4.2.3 Implementation Details
We select two representative pre-training methods for unsupervised (BEV-MAE [28]) and supervised (AD-PT [25]) branches, respectively. For pre-training phase, we adopt commonly used 3D and 2D backbones in [11, 12, 13] and $N_{\text{cls}}=15$, $\lambda=1$. We train 30 epochs with the Adam optimizer, using the one-cycle policy with a learning rate of 0.003. For the downstream detection task, we train 30 epochs for nuScenes, 80 epochs for KITTI and ONCE. For the downstream segmentation task, we train 20 and 10 epochs for SemanticKITTI and nuScenes, respectively. Our experiments are implemented based on 3DTrans [67], using 8 NVIDIA Tesla A100 GPUs. Note that our experiments are under label-efficiency setting, which means that we conduct fine-tuning on a randomly selected subset of the downstream datasets (e.g., $5\%$ for nuScenes detection task, $20\%$ for KITTI and ONCE detection tasks, and $10\%$ for SemanticKITTI and nuScenes segmentation tasks).
@@ 4.3.3 ONCE Detection
Figure 5: Fine-tuning on ONCE dataset for detection task, where 20% training data are used
TABLE IV: Few-shot performance on nuScenes for segmentation task using 100% pre-training data. We fine-tune on 5% and 10% nuScenes training data, respectively, and show the results of some of the categories.
@@ 4.3.4 SemanticKITTI Segmentation
TABLE V: Semi-supervised pre-training setting. For downstream on nuScenes, only 5% training data are used. L5% denotes that we perform pre-training on 5% sequence-level labeled data, while U5% represents 5% unlabeled data.
@@
SPOT (SEMI) L5% + U5% 5% 38.50 50.03
SPOT (SEMI) L5% + U15% 5% 39.81 51.51
@@
SPOT (SEMI) L5% + U5% 5% 43.65 53.82
SPOT (SEMI) L5% + U15% 5% 45.18 54.98
TABLE VI: Semi-supervised pre-training setting. Fine-tune using 20% training data on KITTI for detection task.
@@
SPOT (SEMI) L5% + U5% 20% 65.18
SPOT (SEMI) L5% + U15% 20% 66.45
@@
SPOT (SEMI) L5% + U5% 20% 70.40
SPOT (SEMI) L5% + U15% 20% 70.86
TABLE VII: Semi-supervised pre-training setting. We fine-tune using 5% training data on nuScenes benchmark for segmentation downstream task.
@@
SPOT (SEMI) L5% + U5% 5% 47.60
SPOT (SEMI) L5% + U15% 5% 48.84
TABLE VIII: Semi-supervised pre-training setting. We fine-tune using 10% training data on SemanticKITTI benchmark for segmentation downstream task.
@@
SPOT (SEMI) L5% + U5% 10% 53.62
SPOT (SEMI) L5% + U15% 10% 54.70

### v4

@@ 1 Introduction
To overcome both task-level and dataset-level gaps and learn general representations, we propose SPOT, namely Scalable Pre-training via Occupancy prediction for learning Transferable representation. Our key innovation lies in establishing a unified "one-for-all" pre-training paradigm that enables a single pre-training session to generalize across multiple tasks (detection, segmentation), datasets (Waymo [22], KITTI [18], nuScenes [21], ONCE [20], SemanticKITTI [19]), and sensor configurations, addressing the fundamental limitation of existing task-specific approaches. Firstly, we argue that occupancy prediction serves as a more general pre-training task for task-level generalization, as compared to 3D object detection and LiDAR semantic segmentation. The reason lies in that occupancy prediction is based on denser voxel-level labels with abundant classes, which incorporates spatial information similar to 3D object detection as well as semantic information introduced in semantic segmentation. We provide the first rigorous theoretical foundation through information-theoretic analysis that justifies semantic occupancy prediction as an effective 3D pre-training task. Besides, we consider temporally sufficient representations in the context of 3D pre-training, which contains the information shared among consecutive frames, and theoretically explain why the proposed occupancy-based pre-training outperforms self-supervised methods like MAE in downstream tasks for autonomous driving. Secondly, as the existing datasets use LiDAR sensors with various numbers of laser beams and different category annotation strategies, we propose to use beam re-sampling for point cloud augmentation and class-balancing strategies to overcome these domain gaps. Beam re-sampling augmentation simulates LiDAR sensors with different numbers of laser beams to augment point clouds from a single source pre-training dataset, alleviating the domain gap brought by LiDAR types. Class-balancing strategies apply balance sampling on the dataset and category-specific weights on the loss functions to narrow down the annotation gap. Beyond component integration, SPOT features several engineering insights specifically tailored for 3D pre-training scenarios, including the choice of 2D decoder over traditional 3D decoders which reduces pre-training time from 31 hours to 2.5 hours per epoch while improving parameter efficiency and downstream generalization. Furthermore, our semi-supervised and weakly-supervised pre-training experiments as described in Sec. 4.4 provide strong evidence that SPOT consistently enhances the performance of different downstream tasks, even without using any human-annotated labels. Last but not least, we observe that using larger amounts of pre-training data leads to better performance on various downstream tasks. This holds true even when the pre-training data are generated through pseudo-labeling. These findings indicate that SPOT is a scalable pre-training method for LiDAR point clouds, paving the way for large-scale 3D representation learning in autonomous driving.
@@ 3 The Proposed Method
We discuss the proposed SPOT in detail. As shown in Fig. 2, SPOT contains four parts: (a) Augmentations on LiDAR point clouds. (b) Encoder for LiDAR point clouds to generate BEV features, which are pre-trained and used for different downstream architectures and tasks. (c) Decoder to predict occupancy based on BEV features. (d) Loss function with class-balancing strategy. We first introduce the problem formulation as well as the overall pipeline in Sec. 3.1. Then we respectively discuss beam re-sampling augmentation and class-balancing strategies in Sec. 3.2 and Sec. 3.3. In Sec. 3.4, we provide a theoretical analysis to demonstrate temporally sufficient representations for pre-training in the scenario of autonomous driving.
@@ 3.3.1 Dataset Balancing
where $N_{i}$ is the number of samples for the $i^{th}$ class. Fewer samples in a category brings higher weight $s_{i}$ for it. The square-root rebalancing formula is designed to provide a balanced approach that addresses class imbalance while preventing excessive duplication of rare categories.
To implement this strategy, we perform frame-level random duplication to maintain spatial-temporal consistency. Specifically, for each underrepresented category, we identify scenes containing instances of that category and apply weighted random sampling with replacement based on the computed weights $s_{i}$. When a frame is selected for duplication, all associated annotations are duplicated together to preserve coherence. This frame-level duplication strategy enables more effective learning of general scene representations during pre-training, thereby facilitating improved performance on downstream tasks.
@@ 4 Experiments
The goal of pre-training is to learn general representations for various downstream tasks, datasets, and architectures. We design extensive experiments to answer the question whether SPOT learns such representations in a label-efficiency way. We introduce the employed datasets and experiment setup in Sec. 4.1 and 4.2, respectively, followed by main results with different baselines in Sec. 4.3. Then in Sec. 4.4, we further conduct semi-supervised and weakly-supervised pre-training experiments specifically to demonstrate the applicability of SPOT in the case of utilizing very small part of annotations, and SPOT consistently demonstrates excellent performance on downstream tasks. Finally, we provide discussions about upstream pre-training and downstream fine-tuning, ablation study and visualization results in Sec. 4.5, Sec. 4.6 and Sec. 4.7.
@@ 4.1.4 ONCE Dataset
ONCE [20] is a large-scale autonomous dataset collected in China using a 40-beam LiDAR. It encompasses a diverse range of data collected at various times, under different weather conditions, and across multiple regions. The dataset comprises over one million frames of point cloud data, with approximately 15K frames containing annotations. The remaining unlabeled point cloud data serves as resources for weakly-supervised and semi-supervised algorithms.
TABLE I: Few-shot performance of SPOT on nuScenes validation set. P.D.A. denotes Pre-training Data Amount. We fine-tune on 5% nuScenes training data.
@@
TABLE II: Few-shot performance ($\text{AP}_{\text{3D}}$) of SPOT on KITTI validation set. P.D.A. represents the Pre-training Data Amount, and fine-tuning is performed on 20% KITTI training data.
@@ 4.1.5 SemanticKITTI Dataset
SemanticKITTI dataset [19] is a large-scale dataset based on the KITTI vision, collected by a 64-beam LiDAR sensor. It has 22 sequences, of which sequences 0-7 and 9-10 are used as the training set (19K frames in total), and sequence 8 (4K frames) is used as the validation set, and the remaining 11 sequences (20K frames) as the test set.
@@ 4.2 Experimental Setup
TABLE III: Few-shot performance of SPOT on SemanticKITTI validation set for segmentation task using 100% pre-training data. We fine-tune on 10% training data and show the results of some of the categories.
@@ 4.2.3 Implementation Details
We select two representative pre-training methods for unsupervised (BEV-MAE [28]) and supervised (AD-PT [25]) branches, respectively. For pre-training phase, we adopt commonly used 3D and 2D backbones in [11, 12, 13] and $N_{\text{cls}}=15$, $\lambda=1$. We train 30 epochs with the Adam optimizer, using the one-cycle policy with a learning rate of 0.003. For the downstream detection task, we train 30 epochs for nuScenes, 80 epochs for KITTI and ONCE. For the downstream segmentation task, we train 20 and 10 epochs for SemanticKITTI and nuScenes, respectively. Our experiments are implemented based on 3DTrans [67], using 8 NVIDIA Tesla A100 GPUs. Note that our experiments are under label-efficiency setting, which means that we conduct fine-tuning on a randomly selected subset of the downstream datasets (e.g., $5\%$ for nuScenes detection task, $20\%$ for KITTI and ONCE detection tasks, and $10\%$ for SemanticKITTI and nuScenes segmentation tasks). All evaluation results reported in this paper are conducted on the validation split of the respective datasets.
@@ 4.3.3 ONCE Detection
Figure 5: Fine-tuning on ONCE validation set for detection task, where 20% training data are used
TABLE IV: Few-shot performance on nuScenes validation set for segmentation task using 100% pre-training data. We fine-tune on 5% and 10% nuScenes training data, respectively, and show the results of some of the categories.
@@ 4.3.4 SemanticKITTI Segmentation
TABLE V: Label-efficient pre-training setting, where SEMI and WS denote the semi-supervised and weakly-supervised pre-training setting, respectively. For downstream on nuScenes, only 5% training data are used. L5% denotes that we perform pre-training on 5% sequence-level labeled data, while W5% represents 5% weakly-labeled data.
@@
SPOT L100% 5% 42.57 54.28
SPOT (SEMI) L5% + W5% 5% 38.50 50.03
SPOT (SEMI) L5% + W15% 5% 39.81 51.51
SPOT (SEMI) L5% + W95% 5% 42.18 54.44
SPOT (WS) W100% 5% 42.24 54.35
@@
SPOT L100% 5% 47.47 57.11
SPOT (SEMI) L5% + W5% 5% 43.65 53.82
SPOT (SEMI) L5% + W15% 5% 45.18 54.98
SPOT (SEMI) L5% + W95% 5% 47.58 56.90
SPOT (WS) W100% 5% 47.56 57.18
TABLE VI: Label-efficient pre-training setting. Fine-tuning process uses 20% KITTI training data and we evaluate on KITTI validation set for detection task.
@@
SPOT L100% 20% 67.36
SPOT (SEMI) L5% + W5% 20% 65.18
SPOT (SEMI) L5% + W15% 20% 66.45
SPOT (SEMI) L5% + W95% 20% 67.66
SPOT (WS) W100% 20% 67.68
@@
SPOT L100% 20% 71.77
SPOT (SEMI) L5% + W5% 20% 70.40
SPOT (SEMI) L5% + W15% 20% 70.86
SPOT (SEMI) L5% + W95% 20% 71.67
SPOT (WS) W100% 20% 72.00
TABLE VII: Label-efficient pre-training setting. Fine-tuning process uses 5% nuScenes training data and we evaluate on nuScenes validation set for segmentation downstream task.
@@
SPOT L100% 5% 49.88
SPOT (SEMI) L5% + W5% 5% 47.60
SPOT (SEMI) L5% + W15% 5% 48.84
SPOT (SEMI) L5% + W95% 5% 50.17
SPOT (WS) W100% 5% 51.07
TABLE VIII: Label-efficient pre-training setting. Fine-tuning process uses 10% SemanticKITTI training data and we evaluate on SemanticKITTI validation set for segmentation downstream task.
@@
SPOT L100% 10% 55.58
SPOT (SEMI) L5% + W5% 10% 53.62
SPOT (SEMI) L5% + W15% 10% 54.70
SPOT (SEMI) L5% + W95% 10% 55.96
SPOT (WS) W100% 10% 56.18
4.4 From Supervised to Semi-supervised and Weakly-supervised Pre-training
In this section, considering that SPOT requires supervised information (for the purpose of generating dense occupancy) to perform the 3D pre-training task, we study SPOT’s dependence on pre-training supervision information. In order to demonstrate SPOT’s ability to scale up, we design experiments to explore the semi-supervised and weakly-supervised setting during the pre-training phase.
For semi-supervised pre-training setting, we first pre-train the backbone with SPOT using only 5% sequence-level labeled data and 5%, 15%, 95% sequence-level unlabeled data, where the unlabeled data are pseudo-labeled [68] by employing a naive mean-teacher approach [69]. Refer to [69] for more details in calculating the pseudo-labels.
For weakly-supervised pre-training setting, we first employ vision foundation models [70, 71] to obtain semantic labels for foreground and background objects of the 2D image. Then we establish correspondences between the 2D image space and 3D point space to transfer 2D semantic labels to 3D point cloud data. In our practical implementation, we utilize the 3D projection API function [^1] https://github.com/waymo-research/waymo-open-dataset provided by the Waymo dataset [22], which compensates for point cloud motion, alongside effectively correcting distortion and deformation caused by the camera rolling shutter effect, thereby ensuring precise registration between camera and LiDAR data. Through this projection mapping relationship, we associate 3D point cloud data with semantic labels from the image plane, subsequently generating the occupancy labels required for SPOT pre-training.
After the pre-training phase, the pre-trained backbone is fine-tuned on downstream tasks including nuScenes, KITTI detection tasks, and nuScenes and SemanticKITTI segmentation tasks using different baseline models.
The experimental results of semi-supervised and weakly-supervised pre-training setting are reported in Tables VVIVII, and VIII. It can be found that semi-supervised and weakly-supervised pre-training with SPOT achieves comparable downstream performance as that of fully-supervised pre-training. It consistently improves different architectures on various datasets and tasks. Also, when incorporating more weakly-labeled data to perform the pre-training (e.g., comparing L5%+W5%, L5%+W15% and L5%+W95%), the performance of the downstream task significantly improves. Notably, weakly-supervised pre-training (W100%) also demonstrates competitive performance without using any human-annotated labels during the pre-training phase. Thus, we believe that SPOT is able to generalize to label-efficient pre-training settings and further attain performance scalability on different downstream datasets and tasks such as 3D detection and segmentation tasks.
Overall, we conclude that SPOT requires a certain amount of supervised information in the pre-training dataset, but it remains compatible with unlabeled data (as observed in Table V to VIII). From another perspective, SPOT can alleviate the model’s reliance on human annotations in downstream tasks while achieving better model performance, thereby reducing the annotation costs associated with these downstream tasks.
TABLE IX: Fine-tuing on 100% KITTI training data and evaluating on KITTI validation set with 40 recall positions at moderate difficulty level. SPOT (WS) means we pre-train SPOT without manually labeled data, using generated labels by the completely weakly-supervised method as described in Section 4.4. UN means unsupervised method.
     Backbone      Method      F.D.A.      mAP      Car      Pedestrian      Cyclist
     SECOND [11]      From Scratch      100%      65.35      81.50      48.82      65.72
     Occupancy-MAE (UN) [45]      100%      68.24      81.98      53.67      69.08
     SPOT (WS)      100%      69.73      82.52      56.72      69.95
     PV-RCNN [13]      From Scratch      100%      70.57      84.50      57.06      70.14
     Occupancy-MAE (UN) [45]      100%      73.29      84.82      59.07      75.68
     SPOT (WS)      100%      73.14      84.86      61.29      73.27
TABLE X: Fine-tuing on 100% KITTI training data and evaluating on KITTI validation set with AP calculated by 11 recall positions evaluating bounding box and orientation.
Evaluation Method F.D.A. Car Pedestrian Cyclist
Easy Mod. Hard Easy Mod. Hard Easy Mod. Hard
bbox SECOND [11] 100% 90.73 89.76 88.94 68.70 65.27 62.52 87.88 75.43 71.67
Occupancy-MAE (UN) [45] + SECOND 100% 94.81 89.98 89.35 70.37 67.45 65.14 91.82 78.65 73.77
SPOT (WS) + SECOND 100% 95.48 90.22 89.43 72.94 69.08 66.31 93.64 77.34 74.03
aos SECOND [11] 100% 90.73 89.63 88.70 63.46 60.13 56.93 87.63 74.67 71.00
Occupancy-MAE (UN) [45] + SECOND 100% 94.66 89.88 88.92 65.33 61.55 59.23 91.57 78.42 73.50
SPOT (WS) + SECOND 100% 95.44 90.14 89.26 70.03 65.45 62.13 93.19 76.79 73.51
TABLE XI: Fine-tuing on 100% nuScenes detection training data and evaluating on nuScenes detection validation set.
Backbone Method F.D.A. mAP NDS
CenterPoint [12] From Scratch 100% 56.0 64.5
Occupancy-MAE (UN) [45] 100% 56.5 65.0
SPOT (WS) 100% 57.5 65.3
TABLE XII: Fine-tuing on 100% nuScenes segmentation training data and evaluating on nuScenes segmentation validation set.
Backbone Method F.D.A Epoch mIOU
Cylinder3D [15] From Scratch 100% 15 70.22
Occupancy-MAE (UN) [45] 100% 15 71.61
SPOT (WS) 100% 15 72.37
From Scratch 100% 25 70.83
Occupancy-MAE (UN) [45] 100% 25 72.85
SPOT (WS) 100% 25 73.26
4.5 Comparison with Occupancy-based Pre-training Methods
To validate the effectiveness of our pre-training strategy compared to binary occupancy prediction methods, we conduct extensive experiments on both detection and segmentation downstream tasks. Specifically, we evaluate our method SPOT (WS), which is pre-trained using the auto-generated labels without manually labeled data (see Sec. 4.4 for details on SPOT’s unlabeled pre-training), against the recent binary occupancy prediction pre-training method, Occupancy-MAE [45]. All the experimental settings in this section are consistent with those reported in their paper.
4.5.1 Detection Results
As shown in Tables IX and X, we evaluate the detection performance on the KITTI validation set using SECOND and PV-RCNN as detectors. For SECOND, our method achieves 69.73% mAP on moderate difficulty, outperforming Occupancy-MAE [45] by 1.49% (69.73% vs 68.24%). The improvements are consistent across different categories, with notable gains of 0.54% on Car, 3.05% on Pedestrian, and 0.87% on Cyclist. Similar trends can be observed on PV-RCNN, where our method obtains better performance on Car and Pedestrian categories while maintaining competitive results on Cyclist detection.
We further validate the effectiveness on the more challenging nuScenes dataset. As shown in Table XI, with CenterPoint as the detector, our method achieves 57.5% mAP and 65.3% NDS, surpassing Occupancy-MAE [45] by 1.0% and 0.3% respectively. The consistent improvements across different datasets and backbone networks demonstrate that our semantic occupancy prediction provides richer supervision signals than binary occupancy prediction for learning transferable representations.
4.5.2 Segmentation Results
To evaluate the generalization ability of our pre-training strategy, we also conduct experiments on the segmentation task using Cylinder3D as the backbone. As shown in Table XII, our method consistently outperforms Occupancy-MAE [45] under different training epochs. Specifically, with 15 epochs of fine-tuning, our method achieves 72.37% mIOU, surpassing Occupancy-MAE [45] by 0.76%. When training for longer epochs, the performance gap is maintained (73.26% vs 72.85%). The superior segmentation results further verify that learning to predict semantic occupancy helps the model better understand the 3D scene structure and semantic information, which benefits various downstream tasks.
The consistent improvements compared with Occupancy-MAE [45] across different tasks, backbones and datasets demonstrate the superiority of our SPOT pre-training framework. Unlike Occupancy-MAE that requires dataset-dependent pre-training before fine-tuning on each specific downstream task, our framework advocates a unified "one-to-many" paradigm where a single pre-training on Waymo dataset enables effective transfer to various downstream tasks and datasets. More importantly, even under the same label-free setting, SPOT consistently outperforms Occupancy-MAE by significant margins across multiple tasks (detection, segmentation) and datasets (KITTI, nuScenes), which underscores the effectiveness of our SPOT framework in leveraging semantic-aware occupancy prediction to learn robust, transferable, and domain-agnostic 3D representations.